# Inversion-DPO: Precise and Efficient Post-Training for Diffusion Models

### v2

@@ 5.4 Quantitative Results
SDXL [43] 0.223 0.334 6.07
Diffusion-DPO [57] 0.227 0.340 6.05
DDPO [3] 0.222 0.336 6.10
D3PO [65] 0.223 0.338 5.99
Demon [69] 0.220 0.336 6.15
IterComp [71] 0.232 0.340 6.22
@@
Table 1: Quantitative results of typical text-to-image generation. $\mathcal{y}$ indicates that the model is built upon SD1.5. The same applies to the following tables.
@@
SGDiff [66] 35.8 0.304 0.787 0.698
SG-Adapter [51] 27.8 0.314 0.771 0.693
R3CD [34] 27.0 0.342 0.803 0.715
SDXL-SG [31] 26.7 0.340 0.792 0.703
@@
Table 2: Quantitative results on the more challenging task of compositional image generation. $\ast$ means the results of Inversion-DPO trained on the compositional paired dataset.
Table 1 reports the quantitative performance on typical text-to-image generation for both baseline models and ours. Compared to base T2I models [43, 47], post-training methods [57, 3, 65, 71], and training-free approach [69], Inversion-DPO achieves state-of-the-art results in terms of human preference alignment (PickScore), text-image consistency (CLIP Score), and aesthetic quality (Aesthetic Score). In addition to the formal SDXL-based version of Inversion-DPO, we also validate its adaptability on SD1.5.
We further investigate the performance of the proposed method on compositional image generation. As in Tab. 2, after fine-tuning on the compositional dataset, Inversion-DPO achieves the best image quality as measured by FID, outperforming existing baseline models. Moreover, it obtains the highest scores on complexity evaluation metrics, which comprehensively assess the fidelity of the generated images at the scene, entity, and relation levels.
Empirically, our method also brings notable improvements in training efficiency. In Fig. 2, Inversion-DPO achieves superior performance with significantly fewer training steps compared to Diffusion-DPO. Due to hardware limitation, our reimplementation of Diffusion-DPO with the official implementation does match the performance of the official checkpoint. Besides, we observe a slight decay of performance at the begining which may be a result of the instability of Diffusion-DPO’s training. Nevertheless, under the same resources, our Inversion-DPO constantly improves the performance of the generative model. Specifically, our variants with only 20-step DDIM Inversion reaches a performance of 0.2247 at 1000 steps, comparable to Diffusion-DPO’s 0.2245 at 2000, over 2$\times$ faster. Inversion-DPO with 80 steps is over 4$\times$ faster. Furthermore, at 2000 steps Inversion-DPO with 80-step inversion also outperforms the reported performance of Diffusion-DPO as labeled by the dotted line, further demonstrating the advantages of our method in both precision and efficiency. Results on other metrics are in Supplementary Material.
5.5 Ablation Study
@@
Table 3: Effect of inversion steps on model performance.
Analysis of the different inversion steps. We analyze the impact of different DDIM inversion steps on model training, with results shown in Tab. 3 and Fig. 2. It is evident that as the number of inversion steps increases, the model consistently achieves better performance across all metrics. This can be attributed to the fact that with more DDIM inversion steps, the distributional gap between adjacent steps becomes smaller, satisfying the assumption in Sec. 3.2 and making it easier to guide the model toward correct optimization directions. The configuration with 80 inversion steps is ultimately selected for training Inversion-DPO for a balance between performance and computational cost.
6 Conclusion
@@
S7 User Study
To evaluate the alignment between the generated image and human cognition, we design a user study involving two control groups: the group of SDXL [43] vs. Inversion DPO and the group of Diffusion-DPO [57] vs. Inversion-DPO. For each control group, we randomly select 500 pairs of images. In each trial, users are presented with two images and their corresponding prompts (text descriptions), one pair coming from SDXL and Inversion-DPO, and the other from Diffusion-DPO and Inversion-DPO. Users’ task is to choose the image that best matched the corresponding prompt. It is important to note that some of the generated images contain inappropriate content, so we remove a few images to ensure the experiment adhered to ethical guidelines and to prevent any potential negative impact.
@@
Table S4: Comparison of user preference between different image generation methods. In the comparison, Inversion-DPO consistently shows a higher preference from users, indicating its better alignment with human cognition and image content.
@@
S8 Additional Generation Examples
Figure S5: More high-quality and visually appealing generation examples.
@@
S9 Additional Explanation of Efficiency
Figure S6: Relationship between training steps and Aesthetics Score improvements. The orange curve shows the performance of our Inversion-DPO trained with different numbers of DDIM inversion steps. The gray curve represents our implementation of Diffusion-DPO [57], while the dashed line corresponds to the officially released weights of Diffusion-DPO. All models are trained under identical configurations.
Figure S7: Relationship between training steps and CLIP Score improvements. The red curve shows the performance of our Inversion-DPO trained with different numbers of DDIM inversion steps. The gray curve represents our implementation of Diffusion-DPO [57], while the dashed line corresponds to the officially released weights of Diffusion-DPO. All models are trained under identical configurations.
In addition to the relationship between training steps and PickScore [26] presented in the main paper, we also report the correlations between training steps and Aesthetics Score [49] (Fig. S6), as well as between training steps and CLIP Score [44] (Fig. S7). Under identical training configurations, Inversion-DPO consistently outperforms Diffusion-DPO across all metrics, while also demonstrating faster convergence. Notably, even under constrained training conditions due to hardware limitations, Inversion-DPO still surpasses the officially released Diffusion-DPO in final performance.
S10 Visual Interpretation of the Underlying Assumptions
Figure S8: Visualization of data distributions in the inversion and sampling processes.
In our derivation, a key assumption based on DDIM inversion [52] is that the values of $\boldsymbol{x}_{t-1}^{w}$ are similar when obtained from the inversion of $x_{0}^{w}$ and from the forward sampling of $x_{T}^{w}$. To assess the validity of this assumption, we perform real-time visualization of the distributions along both the inversion and sampling trajectories. As illustrated in Fig. S8, the visual representations at corresponding time steps exhibit high similarity in both the forward and reverse processes, providing intuitive evidence that supports the reliability of this assumption.
S11 Detailed Derivation of InversionDPO
@@
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\left[\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})\right]$ (S18)
@@
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\left[\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})\right]$ (S19)
Since $p_{\theta}$ and $p_{\theta_{0}}$ share the same distribution of $x_{T}$ as standard Gaussian, $\log p_{\theta}(x_{T}^{w})-\log p_{\theta_{0}}(x_{T}^{w})$ cancel out. The conditional probabilities of $p_{\theta}$ and $p_{\theta_{0}}$ are Gaussian distributions as in equation 8 and becomes dirac-delta as $\sigma_{t}$ approaches 00.
@@
$\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\boldsymbol{x}_{t}^{w}+$ (S20)
$\left(\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}-\frac{\sqrt{\alpha_{t-1}({1-\alpha_{t}})}}{\sqrt{\alpha_{t}}}\right)\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)$ (S21)
@@
$-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})$ $=\frac{\|\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})-\boldsymbol{x}_{t-1}^{w}\|}{2\sigma_{t}^{2}}^{2}+C$ (S22)
@@
$\|\boldsymbol{\mu}_{\theta}-\boldsymbol{\mu}_{\theta_{0}}\|^{2}=\frac{1-\alpha_{t}}{\alpha_{t}}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2}$ (S23)
@@
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\frac{1-\alpha_{t}}{2\sigma_{t}^{2}\alpha_{t}}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2},$ (S24)
We adopt the simple form as in DDPM [17] to eliminate the weighs and coefficients. Therefore, taking the winning and lossing sample all together, we reach the final loss as
$\mathcal{L}_{\text{Inversion-DPO}}(\theta)=-\mathbb{E}_{(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l})}\log\sigma\bigg{(}\beta\mathbb{E}_{\begin{subarray}{c}\boldsymbol{x}_{1:T}^{w}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{w}|\boldsymbol{x}_{0}^{w})\\ \boldsymbol{x}_{1:T}^{l}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{l}|\boldsymbol{x}_{0}^{l})\end{subarray}}$ (S25)
@@
S12 Impact of the imprecise approximation
@@
Table S5: Performance impact of approximating the reverse distribution $p_{\theta}$ with the forward $q$. In Inversion-DPO, the true reverse distribution $p_{\theta}$ is utilized, whereas Diffusion-DPO relies on its approximation $q$.
In Diffusion-DPO, due to the inaccessibility of the reverse trajectory, the posterior sampling $p_{\theta}(x_{1:T}|x_{0})$ is approximated by the noisy process $q(x_{1:T}|x_{0})$. We argue that this approximation introduces non-negligible training errors and conduct an ablation study to investigate this effect. While Inversion-DPO involves recovering the whole noise trajectory across multiple time steps for alignment, in the code implementation, the original Diffusion-DPO samples a single timestep from $q(x_{1:T}|x_{0})$ and augment with a value $T$ to obtain the final result. Such difference may also introduce performance discrepancy. For a fair comparison, we also evaluate Inversion-DPO by randomly sampling a single timestep from $p_{\theta}(x_{1:T}|x_{0})$. As presented in Tab. S5 in SDXL and SD1.5, training with DDIM Inversion with $p_{\theta}$ consistently yields better results than approximation based on the forward distribution $q$, highlighting the necessity and effectiveness of the core insight behind Inversion-DPO.
S13 A Brief Introduction of Stable Diffusion
Diffusion models, such as Denoising Diffusion Probabilistic Models (DDPMs) [17] and Denoising Diffusion Implicit Models (DDIMs) [52], provide a generative modeling framework that is based on the gradual corruption of data through a noise injection process followed by the recovery of the original data. These models leverage a Markov chain process to progressively add Gaussian noise to the data in a forward process, and then learn to reverse this process to recover the data from noisy samples.
@@
To accelerate the sampling process, [52] introduces DDIMs, which extend DDPMs by generalizing the forward process to be non-Markovian. This modification allows for a more efficient sampling process that reduces the number of steps required to generate high-quality samples. In DDIMs, the generative process is deterministic, and the model can sample from fewer timesteps while maintaining high sample quality. The key insight behind DDIMs is that the reverse process in DDPM can be reparameterized using a deterministic trajectory, which simplifies the sampling process significantly.
S14 Limitation
@@
Figure S9: Additional results of failure examples.

### v4

@@ 5.4. Quantitative Results
SDXL (Podell et al., 2023) 0.223 0.334 6.07
Diffusion-DPO (Wallace et al., 2024) 0.227 0.340 6.05
DDPO (Black et al., 2024) 0.222 0.336 6.10
D3PO (Yang et al., 2024b) 0.223 0.338 5.99
Demon (Yeh et al., 2025) 0.220 0.336 6.15
IterComp (Zhang et al., 2024c) 0.232 0.340 6.22
@@
Table 1. Quantitative results of typical text-to-image generation. $\mathcal{y}$ indicates that the model is built upon SD1.5. The same applies to the following tables.
@@
SGDiff (Yang et al., 2022) 35.8 0.304 0.787 0.698
SG-Adapter (Shen et al., 2024) 27.8 0.314 0.771 0.693
R3CD (Liu and Liu, 2024) 27.0 0.342 0.803 0.715
SDXL-SG (Li et al., 2024b) 26.7 0.340 0.792 0.703
@@
Table 2. Quantitative results on the more challenging task of compositional image generation. $\ast$ means the results of Inversion-DPO trained on the compositional paired dataset.
Table 1 reports the quantitative performance on typical text-to-image generation for both baseline models and ours. Compared to base T2I models (Podell et al., 2023; Rombach et al., 2022), post-training methods (Wallace et al., 2024; Black et al., 2024; Yang et al., 2024b; Zhang et al., 2024c), and training-free approach (Yeh et al., 2025), Inversion-DPO achieves state-of-the-art results in terms of human preference alignment (PickScore), text-image consistency (CLIP Score), and aesthetic quality (Aesthetic Score). In addition to the formal SDXL-based version of Inversion-DPO, we also validate its adaptability on SD1.5.
We further investigate the performance of the proposed method on compositional image generation. As in table 2, after fine-tuning on the compositional dataset, Inversion-DPO achieves the best image quality as measured by FID, outperforming existing baseline models. Moreover, it obtains the highest scores on complexity evaluation metrics, which comprehensively assess the fidelity of the generated images at the scene, entity, and relation levels.
Empirically, our method also brings notable improvements in training efficiency. In fig. 3 and fig. 4, Inversion-DPO achieves superior performance with significantly fewer training steps compared to Diffusion-DPO. Due to hardware limitation, our reimplementation of Diffusion-DPO with the official implementation does match the performance of the official checkpoint. Besides, we observe a slight decay of performance at the begining which may be a result of the instability of Diffusion-DPO’s training. Nevertheless, under the same resources, our Inversion-DPO constantly improves the performance of the generative model. Specifically, our variants with only 20-step DDIM Inversion reaches a performance of 0.2247 at 1000 steps, comparable to Diffusion-DPO’s 0.2245 at 2000, over 2$\times$ faster. Inversion-DPO with 80 steps is over 4$\times$ faster. Furthermore, at 2000 steps Inversion-DPO with 80-step inversion also outperforms the reported performance of Diffusion-DPO as labeled by the dotted line, further demonstrating the advantages of our method in both precision and efficiency. Results on other metrics are in Supplementary Material.
5.5. Ablation Study
@@
Table 3. Effect of inversion steps on model performance.
Analysis of the different inversion steps. We analyze the impact of different DDIM inversion steps on model training, with results shown in table 3 and fig. 3. It is evident that as the number of inversion steps increases, the model consistently achieves better performance across all metrics. This can be attributed to the fact that with more DDIM inversion steps, the distributional gap between adjacent steps becomes smaller, satisfying the assumption in Sec. 3.2 and making it easier to guide the model toward correct optimization directions. The configuration with 80 inversion steps is ultimately selected for training Inversion-DPO for a balance between performance and computational cost.
6. Conclusion
@@
7. Acknowledgement
This paper is supported by Provincial Key Research and Development Plan of Zhejiang Province under No. 2024C01250(SD2) and National Natural Science Foundation of China (Grant No. 62006208).
@@ 7. Acknowledgement
Appendix S1 User Study
To evaluate the alignment between the generated image and human cognition, we design a user study involving two control groups: the group of SDXL (Podell et al., 2023) vs. Inversion DPO and the group of Diffusion-DPO (Wallace et al., 2024) vs. Inversion-DPO. For each control group, we randomly select 500 pairs of images. In each trial, users are presented with two images and their corresponding prompts (text descriptions), one pair coming from SDXL and Inversion-DPO, and the other from Diffusion-DPO and Inversion-DPO. Users’ task is to choose the image that best matched the corresponding prompt. It is important to note that some of the generated images contain inappropriate content, so we remove a few images to ensure the experiment adhered to ethical guidelines and to prevent any potential negative impact.
@@
Table S1. Comparison of user preference between different image generation methods. In the comparison, Inversion-DPO consistently shows a higher preference from users, indicating its better alignment with human cognition and image content.
@@
Appendix S2 Additional Generation Examples
Figure S1. More high-quality and visually appealing generation examples.
Figure S2. Qualitative results on compositional image generation. Inversion-DPO proves effective in addressing key challenges in compositional image generation, including object omission (orange and green annotations in (a), (d), and (e)), attribute errors (red annotations in (b), (c), and (d)), and relation inaccuracies (blue annotations in (c) and (d)). As shown in the last row, Inversion-DPO achieves satisfactory generation results on cases where baseline models struggle.
@@
Appendix S3 Additional Explanation of Efficiency
Figure S3. Relationship between training steps and CLIP Score improvements. The red curve shows the performance of our Inversion-DPO trained with different numbers of DDIM inversion steps. The gray curve represents our implementation of Diffusion-DPO (Wallace et al., 2024), while the dashed line corresponds to the officially released weights of Diffusion-DPO. All models are trained under identical configurations.
In addition to the relationship between training steps and PickScore (Kirstain et al., 2023), as well as between training steps and Aesthetics Score (Schuhmann et al., 2022) presented in the main paper, we also report the correlations between training steps and CLIP Score (Radford et al., 2021) (fig. S3). Under identical training configurations, Inversion-DPO consistently outperforms Diffusion-DPO across all metrics, while also demonstrating faster convergence. Notably, even under constrained training conditions due to hardware limitations, Inversion-DPO still surpasses the officially released Diffusion-DPO in final performance.
Appendix S4 Visual Interpretation of the Underlying Assumptions
Figure S4. Visualization of data distributions in the inversion and sampling processes.
In our derivation, a key assumption based on DDIM inversion (Song et al., 2020) is that the values of $\boldsymbol{x}_{t-1}^{w}$ are similar when obtained from the inversion of $x_{0}^{w}$ and from the forward sampling of $x_{T}^{w}$. To assess the validity of this assumption, we perform real-time visualization of the distributions along both the inversion and sampling trajectories. As illustrated in fig. S4, the visual representations at corresponding time steps exhibit high similarity in both the forward and reverse processes, providing intuitive evidence that supports the reliability of this assumption.
Appendix S5 Detailed Derivation of Inversion-DPO
@@
(S1) $\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\left[\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})\right]$
@@
(S2) $\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\left[\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})\right]$
Since $p_{\theta}$ and $p_{\theta_{0}}$ share the same distribution of $x_{T}$ as standard Gaussian, $\log p_{\theta}(x_{T}^{w})-\log p_{\theta_{0}}(x_{T}^{w})$ cancel out. The conditional probabilities of $p_{\theta}$ and $p_{\theta_{0}}$ are Gaussian distributions as in equation 8 and becomes dirac-delta as $\sigma_{t}$ approaches 0.
@@
(S3) $\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\boldsymbol{x}_{t}^{w}+$
(S4) $\left(\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}-\frac{\sqrt{\alpha_{t-1}({1-\alpha_{t}})}}{\sqrt{\alpha_{t}}}\right)\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)$
@@
(S5) $-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})$ $=\frac{\|\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})-\boldsymbol{x}_{t-1}^{w}\|}{2\sigma_{t}^{2}}^{2}+C$
@@
(S6) $\|\boldsymbol{\mu}_{\theta}-\boldsymbol{\mu}_{\theta_{0}}\|^{2}=\frac{1-\alpha_{t}}{\alpha_{t}}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2}$
@@
(S7) $\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\frac{1-\alpha_{t}}{2\sigma_{t}^{2}\alpha_{t}}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2},$
We adopt the simple form as in DDPM (Ho et al., 2020) to eliminate the weighs and coefficients. Therefore, taking the winning and lossing sample all together, we reach the final loss as
(S8) $\mathcal{L}_{\text{Inversion-DPO}}(\theta)=-\mathbb{E}_{(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l})}\log\sigma\bigg{(}\beta\mathbb{E}_{\begin{subarray}{c}\boldsymbol{x}_{1:T}^{w}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{w}|\boldsymbol{x}_{0}^{w})\\ \boldsymbol{x}_{1:T}^{l}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{l}|\boldsymbol{x}_{0}^{l})\end{subarray}}$
@@
Appendix S6 Impact of the Imprecise Approximation
@@
Table S2. Performance impact of approximating the reverse distribution $p_{\theta}$ with the forward $q$. In Inversion-DPO, the true reverse distribution $p_{\theta}$ is utilized, whereas Diffusion-DPO relies on its approximation $q$.
In Diffusion-DPO, due to the inaccessibility of the reverse trajectory, the posterior sampling $p_{\theta}(x_{1:T}|x_{0})$ is approximated by the noisy process $q(x_{1:T}|x_{0})$. We argue that this approximation introduces non-negligible training errors and conduct an ablation study to investigate this effect. While Inversion-DPO involves recovering the whole noise trajectory across multiple time steps for alignment, in the code implementation, the original Diffusion-DPO samples a single timestep from $q(x_{1:T}|x_{0})$ and augment with a value $T$ to obtain the final result. Such difference may also introduce performance discrepancy. For a fair comparison, we also evaluate Inversion-DPO by randomly sampling a single timestep from $p_{\theta}(x_{1:T}|x_{0})$. As presented in table S2 in SDXL and SD1.5, training with DDIM Inversion with $p_{\theta}$ consistently yields better results than approximation based on the forward distribution $q$, highlighting the necessity and effectiveness of the core insight behind Inversion-DPO.
Appendix S7 A Brief Introduction of Stable Diffusion
Diffusion models, such as Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) and Denoising Diffusion Implicit Models (DDIMs) (Song et al., 2020), provide a generative modeling framework that is based on the gradual corruption of data through a noise injection process followed by the recovery of the original data. These models leverage a Markov chain process to progressively add Gaussian noise to the data in a forward process, and then learn to reverse this process to recover the data from noisy samples.
@@
To accelerate the sampling process, (Song et al., 2020) introduces DDIMs, which extend DDPMs by generalizing the forward process to be non-Markovian. This modification allows for a more efficient sampling process that reduces the number of steps required to generate high-quality samples. In DDIMs, the generative process is deterministic, and the model can sample from fewer timesteps while maintaining high sample quality. The key insight behind DDIMs is that the reverse process in DDPM can be reparameterized using a deterministic trajectory, which simplifies the sampling process significantly.
Appendix S8 Limitation
@@
Figure S5. Additional results of failure examples.